OpenStudy (anonymous):

Factor Completely x 4 - 19x 2 + 48

OpenStudy (callisto):

x^4 - 19x^2 + 48 = x^4 - 3x^2 - 16x^2 + 48 = x^2 (x^2 -3) - 16 (x^2 -3) = (x^2 -16) ( x^2 -3) Now, use the identity a^2-b^2 = (a-b)(a+b) to factorise (x^2-16) Can you try it?

OpenStudy (anonymous):

I'm so lost here

OpenStudy (anonymous):

wait. I get it. Thank you.

OpenStudy (callisto):

Hmm... what did you get?!